Consider the following two reactions and enthalpies:
1) 2C(s) + O2(g) → 2CO(g) ΔH = – 220 kJ/mol
2) C(s) + O2(g) → CO2(g) ΔH = – 393 kJ/mol
From these two reactions, calculate the heat effect (enthalpy change ΔH) for the reaction:
3) CO2(g) → CO(g) + ½O2(g)
1
Expert's answer
2016-11-22T13:01:14-0500
According to the Hess's law:

ΔH = ½*ΔH 1 – ΔH 2 = -110 + 393 = 283 (kJ/mol)